i need more than 1 problem please. heres the problems:
evaluate: 3a + (k + 5b) - 9=? if a=7 & b=4 & k=10
find the 7th term in the sequence: 20 1/2, 23 1/3, 26 1/6 29...
evaluate: -6a-4b= ? if a=5 and b=3??
please! thanks

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

evaluate: 3a + (k + 5b) - 9=? if a=7 & b=4 & k=10
3*7 + (10 + 5*4)
21 + (10+20)
21 + 30
51
----------------------------

find the 7th term in the sequence: 20 1/2, 23 1/3, 26 1/6 29...
----------------------------

evaluate: -6a-4b= ? if a=5 and b=3
-6*5 - 4*3
-30 - 12
-42
--------------------------------------
Cheers,
Stan H.

Answer by richwmiller(17219)

The rules are very clear one problem at a time not two not three.
Not only that but the rules also say no similar problems.
Why do you think you can ignore the rules? Because you put a whining please?
Also limit yourself to 4 problems sent separately per day.
Believe it or not because you hogged Stan's time for three problems, two people with real problems will not get their problems answered. Is that fair?